# MS-TP-25-28

### v2

@@ VI.2 Pure mass coupling
of the longitudinal modes (the green dashed–dotted and purple dotted lines, respectively) during the last 30 $e$-folds of inflation computed by using the GEF system of equations (93)–(IV.2) truncated at the order $2p_{\rm max}+1=69$. The lower panel of this figure shows the relative error of the GEF solution compared to the numerical result obtained by treating separate Fourier modes in the momentum space and solving the system of Eqs. (66)–(67) with boundary conditions in Eqs. (68)–(69). We will refer to the latter solution as the mode-by-mode (MbM) solution and define the relative error as
@@
Note that the relative error is typically of order a few percent and only takes somewhat larger values, or order 10 % in the vicinity of a sharp drop in the energy density curves shown in the upper panel of Fig. 1. The latter drop occurs in a region where the horizon-crossing momentum $k_{h,L}$ stops growing because of the non-monotonic behavior of the quantity $aH|R|$. Apart from the near-drop regions, the accuracy of the GEF method is very good and the method can be trustfully used for the studies of the vector DM production.
Figure 1: Upper panel: the dependence of the energy densities $\rho_{\mathcal{Q}}$ (green dashed-dotted line) and $\rho_{\mathcal{F}}$ (purple dotted line) on the number of $e$-folds counted from the end of inflation in the model of mass coupling (114) of the vector field to inflaton with $\beta=12$, obtained by using the GEF truncated at the order $2p_{\rm max}+1=69$. Lower panel: evolution of the relative error of the GEF solution with respect to the mode-by-mode solution in momentum space.
Figure 2: The dependence of the energy densities $\rho_{\mathcal{Q}}$ (green dashed-dotted line) and $\rho_{\mathcal{F}}$ (purple dotted line) of the longitudinal modes of the vector field as well as the inflaton energy density $\rho_{\rm inf}$ (black dashed line) on the number of $e$-folds in the model of mass coupling (114) of the vector field to inflaton with $\beta=16$, obtained by using the GEF truncated at the order $2p_{\rm max}+1=63$. The point $N=0$ is chosen at the moment of time when inflation would have ended in the absence of gauge fields. The backreaction prolongs the accelerated expansion of the universe by a few $e$-folds.
Next, in Fig. 2, we show the numerical results for the case of a stronger coupling, $\beta=16$. Here we also plot the inflaton energy density $\rho_{\rm inf}=\dot{\phi}^{2}/2+V(\phi)$ by the thin black dashed line. The produced vector field appears to be strong enough to change the background inflationary evolution and extend the accelerated expansion of the universe by a few $e$-folds. Indeed, the point $N=0$ on the horizontal axis is chosen at the moment of time when inflation would have ended in the absence of gauge fields. As one can see from the figure, inflation still lasts a few $e$-folds more after that if the presence of the vector field is taken into account. In the backreaction regime, the gauge-field energy density stops growing and reveals damped oscillatory behavior which is very analogous to the backreaction regime for the massless gauge fields with the kinetic and axial couplings to the inflaton reported, e.g., in Refs. [57, 74, 58, 65].
@@ VI.3 Kinetic and mass couplings
For the positive value of the coupling constant, $\beta=8$, the electric transverse modes clearly dominate among all gauge-field contributions. The longitudinal modes are subdominant with $\rho_{\mathcal{Q}}$ being slightly larger than $\rho_{\mathcal{F}}$ (the same hierarchy as in the pure mass coupling for $\beta=12$, see Fig. 1). Here again, the error for the dominant energy density, $\rho_{\mathcal{E}}$, remains small, mostly of order 1 %, only reaching the values $\sim 10\,\%$ in a close vicinity of the sharp drop in the energy density. Its origin, similarly to the previous discussion in the case of pure mass coupling, is in the abrupt freezing of the $k_{h}$ evolution in the region when $aHr$ has a non-monotonic behavior. Note that for longitudinal modes the sharp drop is also present (related to $k_{k,L}$ freezing in this case) and happens approximately 1 $e$-folding earlier. The subdominant contributions are determined by GEF with a slightly larger error, from a few to 10 %, and reaching up to 30 % error in the vicinity of the peaks. We would like to emphasize that this increase in the error is absolutely expected in this region as we are comparing two sharply peaked solutions (the ones from the GEF and MbM approaches). Indeed, any tiny mismatch between them leads to a large error as both functions sharply drop by two orders of magnitude near the minimum. In the case of monotonic functions, e.g., for a negative coupling constant $\beta=-6$ shown in Fig. 4, the error of the GEF solution remains well below 10 % for all components of the energy density, including the subleading ones, while for the dominant contribution always remains at the level of 1 %.
Figure 3: Upper panel: the dependence of the energy densities $\rho_{\mathcal{E}}$ (blue solid line), $\rho_{\mathcal{B}}$ (red dashed line), $\rho_{\mathcal{Q}}$ (green dashed-dotted line), and $\rho_{\mathcal{F}}$ (purple dotted line) on the number of $e$-folds counted from the end of inflation in the model of kinetic (113) and mass couplings (114) of the vector field to inflaton with $\beta=8$, obtained by using the GEF truncated at the orders $n_{\rm max}=53$ for the transverse modes and $2p_{\rm max}+1=53$ for the longitudinal modes. Lower panel: evolution of the relative error of the GEF solution with respect to the mode-by-mode solution in momentum space.
Figure 4: The same quantities as shown in Fig. 3 for the case of a negative coupling constant $\beta=-6$. The GEF system was truncated at the orders $n_{\rm max}=71$ for the transverse modes and $2p_{\rm max}+1=71$ for the longitudinal modes.
@@
Now, we switch to the cases with stronger couplings. The results for $\beta=10.4$ are shown in Fig. 5 while for $\beta=-7$ in Fig. 6. Here, the hierarchy between different components of the energy density is the same as in the corresponding cases with weaker couplings. However, for stronger couplings, the produced fields are able to significantly impact the inflationary dynamics. Again, as in the pure mass coupling case, we see that the energy density of the vector field becomes comparable to that of the inflaton field. The time evolution of energy-density components reveals damped oscillations which probably result from the retardation between the changes in the inflaton velocity and the response of the gauge-field subsystem.
The accelerated inflationary expansion lasts longer in the presence of the gauge field. Despite the fact that, for the low-mass vector field, its energy density must redshift as radiation, the energy transfer from the inflaton through the mass and kinetic couplings maintains the gauge-field energy density almost constant and, thus, realizes a vacuum-like state composed from the inflaton and the gauge fields. This extends the duration of inflation by a few $e$-folds.
@@ VII Conclusions
Figure 5: The evolution of the different components of the energy density of the vector field: $\rho_{\mathcal{E}}$ (blue solid line), $\rho_{\mathcal{B}}$ (red dashed line), $\rho_{\mathcal{Q}}$ (green dashed-dotted line), $\rho_{\mathcal{F}}$ (purple dotted line), and the inflaton energy density $\rho_{\rm inf}$ (thin black dashed line) in the model of kinetic (113) and mass couplings (114) of the vector field to inflaton with $\beta=10.4$, obtained by using the GEF truncated at the orders $n_{\rm max}=87$ for the transverse modes and $2p_{\rm max}+1=87$ for the longitudinal modes. The point $N=0$ is chosen at the moment of time when inflation would have ended in the absence of gauge fields.
Figure 6: The same quantities as shown in Fig. 5 for the case of a negative coupling constant $\beta=-7$. The GEF system was truncated at the orders $n_{\rm max}=181$ for the transverse modes and $2p_{\rm max}+1=181$ for the longitudinal modes.
We extended the gradient-expansion formalism (GEF) — previously developed for describing massless Abelian gauge field generation during inflation — to include the longitudinal polarization of a massive vector field. To this end, we derived a coupled system of equations of motion for bilinear functions of the vector field and examined the nonlinear dynamics of massive vector field production, including backreaction on the inflationary background. The evolution of both transverse and longitudinal polarization modes was analyzed in Fourier and position space, with the GEF in the latter case supplemented by appropriate boundary terms.
@@
In the cases without backreaction, we performed computations by employing two methods: the GEF and by treating separate Fourier modes in momentum space. By comparing the results of two approaches, we estimated the relative error of the GEF result. It appears to be of order 1 % for the dominant component of the energy density and a few percent for subleading contributions if all energy densities evolve monotonically. If, on the other hand, sharp peaks or drops are present in the evolution of the energy-density components, the error in these regions may increase to $\sim 20-30\,\%$ for subleading components. Still, this result is fully acceptable for numerical study of the vector DM production during inflation.
@@ Appendix A Equation of motion for longitudinal modes
Introducing a canonically normalized mode function $\psi_{L}$ via Eq. (70), we obtain Eq. (71) for with the coefficient
@@
For the trivial kinetic coupling, $I_{1}\equiv 1$, this expression coincides with the one reported in Refs. [48, 49]. In the limit $\mu\ll 1$ and for $k\gtrsim aH$, the expression simplifies to the one in Eq. (72). Far inside the horizon, equation of motion for $\psi_{L}$ in conformal time $\eta$ takes the oscillator-like form $\psi^{\prime\prime}_{L}(\eta,k)+a^{2}Q\psi_{L}(\eta,k)=0$. To obtain a positive-frequency solution to this equation, we impose the Bunch–Davies boundary condition (79). Using Eq. (70), we translate it to the boundary condition (68) for $\mathcal{A}_{L}$. Further, using Eq. (66), we derive the boundary condition (69) for $\mathcal{D}_{L}$.
@@
Acknowledgements.
S. I. V. is grateful to Professor Oleg Ruchayskiy for his support, useful discussions, and kind hospitality at the Niels Bohr Institute, University of Copenhagen, Denmark, where the final part of this work was done. S. I. V. is supported by the Erasmus+ Mobility grant KA171-2023. The work of A. V. L. was supported by the Swiss National Science Foundation through the Ukrainian–Swiss Joint research project “Cosmic waltz of baryonic and ultralight dark matter: interaction and dynamical interplay” (grant No. IZURZ2_224972). The work of O. O. S. was sustained through the SAFE — Supporting At-Risk Researchers with Fellowships in Europe project, which is funded by the European Union. Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union, the European Research Executive Agency (REA). Neither the European Union nor the granting authority can be held responsible for them.

### v3

@@ VI.2 Pure mass coupling
of the longitudinal modes (the green dashed–dotted and purple dotted lines, respectively) during the last 30 $e$-folds of inflation computed by using the GEF system of equations (93)–(IV.2) truncated at the order $2p_{\rm max}+1=67$. The lower panel of this figure shows the relative error of the GEF solution compared to the numerical result obtained by treating separate Fourier modes in the momentum space and solving the system of Eqs. (66)–(67) with boundary conditions in Eqs. (68)–(69). We will refer to the latter solution as the mode-by-mode (MbM) solution and define the relative error as
@@
Note that the relative error is typically of order a few percent and only takes somewhat larger values, or order 10 % in the vicinity of a sharp drop in the energy density curves shown in the upper panel of Fig. 1. The latter drop occurs in a region where the horizon-crossing momentum $k_{h,L}$ stops growing because of the nonmonotonic behavior of the quantity $aH|R|$. Apart from the near-drop regions, the accuracy of the GEF method is very good and the method can be trustfully used for the studies of the vector DM production.
Figure 2: Dependence of the energy densities $\rho_{\mathcal{Q}}$ (green dashed-dotted line) and $\rho_{\mathcal{F}}$ (purple dotted line) of the longitudinal modes of the vector field as well as the inflaton energy density $\rho_{\rm inf}$ (black dashed line) on the number of $e$-folds in the model of mass coupling (114) of the vector field to inflaton with $\beta=16$, obtained by using the GEF truncated at the order $2p_{\rm max}+1=67$. The point $N=0$ is chosen at the moment of time when inflation would have ended in the absence of gauge fields. The backreaction prolongs the accelerated expansion of the Universe by a few $e$-folds.
Next, in Fig. 2, we show the numerical results for the case of a stronger coupling, $\beta=16$. Here we also plot the inflaton energy density $\rho_{\rm inf}=\dot{\phi}^{2}/2+V(\phi)$ by the thin black dashed line. The produced vector field appears to be strong enough to change the background inflationary evolution and extend the accelerated expansion of the Universe by a few $e$-folds. Indeed, the point $N=0$ on the horizontal axis is chosen at the moment of time when inflation would have ended in the absence of gauge fields. As one can see from the figure, inflation still lasts a few $e$-folds more after that if the presence of the vector field is taken into account. In the backreaction regime, the gauge-field energy density stops growing and reveals damped oscillatory behavior, which is very analogous to the backreaction regime for the massless gauge fields with the kinetic and axial couplings to the inflaton reported, e.g., in Refs. [57, 79, 58, 65].
Figure 3: Upper panel: the dependence of the energy densities $\rho_{\mathcal{E}}$ (blue solid line), $\rho_{\mathcal{B}}$ (red dashed line), $\rho_{\mathcal{Q}}$ (green dashed-dotted line), and $\rho_{\mathcal{F}}$ (purple dotted line) on the number of $e$-folds counted from the end of inflation in the model of kinetic (113) and mass couplings (114) of the vector field to inflaton with $\beta=8$, obtained by using the GEF truncated at the orders $n_{\rm max}=41$ for the transverse modes and $2p_{\rm max}+1=41$ for the longitudinal modes. Lower panel: evolution of the relative error of the GEF solution with respect to the mode-by-mode solution in momentum space.
Figure 4: Same quantities as shown in Fig. 3 for the case of a negative coupling constant $\beta=-6$. The GEF system was truncated at the orders $n_{\rm max}=31$ for the transverse modes and $2p_{\rm max}+1=31$ for the longitudinal modes.
@@ VI.3 Kinetic and mass couplings
For the positive value of the coupling constant, $\beta=8$, the electric transverse modes clearly dominate among all gauge-field contributions. The longitudinal modes are subdominant with $\rho_{\mathcal{Q}}$ being slightly larger than $\rho_{\mathcal{F}}$ (the same hierarchy as in the pure mass coupling for $\beta=12$, see Fig. 1). Here again, the error for the dominant energy density, $\rho_{\mathcal{E}}$, remains small, mostly of order 1 %, only reaching the values $\sim 10\,\%$ in a close vicinity of the sharp drop in the energy density. Its origin, similarly to the previous discussion in the case of pure mass coupling, is in the abrupt freezing of the $k_{h}$ evolution in the region when $aHr$ has a nonmonotonic behavior. Note that for longitudinal modes the sharp drop is also present (related to $k_{k,L}$ freezing in this case) and happens approximately 1 $e$-folding earlier. The subdominant contributions are determined by GEF with a slightly larger error, from a few to 10 %, and reaching up to 30 % error in the vicinity of the peaks. We would like to emphasize that this increase in the error is absolutely expected in this region as we are comparing two sharply peaked solutions (the ones from the GEF and MbM approaches). Indeed, any tiny mismatch between them leads to a large error as both functions sharply drop by two orders of magnitude near the minimum. In the case of monotonic functions, e.g., for a negative coupling constant $\beta=-6$ shown in Fig. 4, the error of the GEF solution remains well below 10 % for all components of the energy density, including the subleading ones, while that for the dominant contribution always remains at the level of 1 %.
Figure 5: The evolution of the different components of the energy density of the vector field: $\rho_{\mathcal{E}}$ (blue solid line), $\rho_{\mathcal{B}}$ (red dashed line), $\rho_{\mathcal{Q}}$ (green dashed-dotted line), $\rho_{\mathcal{F}}$ (purple dotted line), and the inflaton energy density $\rho_{\rm inf}$ (thin black dashed line) in the model of kinetic (113) and mass couplings (114) of the vector field to inflaton with $\beta=10.4$, obtained by using the GEF truncated at the orders $n_{\rm max}=63$ for the transverse modes and $2p_{\rm max}+1=63$ for the longitudinal modes. The point $N=0$ is chosen at the moment of time when inflation would have ended in the absence of gauge fields.
Figure 6: Same quantities as shown in Fig. 5 for the case of a negative coupling constant $\beta=-7$. The GEF system was truncated at the orders $n_{\rm max}=171$ for the transverse modes and $2p_{\rm max}+1=171$ for the longitudinal modes.
@@
Now, we switch to the cases with stronger couplings. The results for $\beta=10.4$ are shown in Fig. 5 while those for $\beta=-7$ are shown in Fig. 6. Here, the hierarchy between different components of the energy density is the same as in the corresponding cases with weaker couplings. However, for stronger couplings, the produced fields are able to significantly impact the inflationary dynamics. Again, as in the pure mass coupling case, we see that the energy density of the vector field becomes comparable to that of the inflaton field. The time evolution of energy-density components reveals damped oscillations that probably result from the retardation between the changes in the inflaton velocity and the response of the gauge-field subsystem.
The accelerated inflationary expansion lasts longer in the presence of the gauge field. Despite the fact that, for the low-mass vector field, its energy density must redshift as radiation, the energy transfer from the inflaton through the mass and kinetic couplings maintains the gauge-field energy density almost constant and, thus, realizes a vacuumlike state composed from the inflaton and the gauge fields. This extends the duration of inflation by a few $e$-folds.
@@ VII Conclusions
We extended the gradient-expansion formalism (GEF) — previously developed for describing massless Abelian gauge field generation during inflation — to include the longitudinal polarization of a massive vector field. To this end, we derived a coupled system of equations of motion for bilinear functions of the vector field and examined the nonlinear dynamics of massive vector field production, including backreaction on the inflationary background. The GEF method is fully equivalent to the conventional mode-by-mode treatment in momentum space. The main difference is technical: Instead of evolving separate Fourier modes (which all become coupled in the nonlinear regime of evolution), the GEF operates directly in position space and evolves a set of bilinear gauge-field correlators that automatically include the contributions of all Fourier modes simultaneously. This makes the method considerably faster and computationally less expensive, especially when the gauge-field production is strong and many modes become coupled in a nonlinear way. The evolution of both transverse and longitudinal polarization modes was analyzed with the GEF and the result was cross-checked by means of the mode-by-mode approach in Fourier space.
@@
In the cases without backreaction, we performed computations by employing two methods: by the GEF and by treating separate Fourier modes in momentum space. By comparing the results of two approaches, we estimated the relative error of the GEF result. It appears to be of order 1 % for the dominant component of the energy density and a few percent for subleading contributions if all energy densities evolve monotonically. If, on the other hand, sharp peaks or drops are present in the evolution of the energy-density components, the error in these regions may increase to $\sim 20-30\,\%$ for subleading components. Still, this result is fully acceptable for numerical study of the vector DM production during inflation.
@@
Acknowledgements.
S. I. V. is grateful to Professor Oleg Ruchayskiy for his support, useful discussions, and kind hospitality at the Niels Bohr Institute, University of Copenhagen, Denmark, where the final part of this work was done. S. I. V. is supported by the Erasmus+ Mobility Grant No. KA171-2023. The work of A. V. L. was supported by the Swiss National Science Foundation through the Ukrainian–Swiss Joint research project “Cosmic waltz of baryonic and ultralight dark matter: interaction and dynamical interplay” (Grant No. IZURZ2_224972). The work of O. S. has received funding through the SAFE — Supporting At-Risk Researchers with Fellowships in Europe project, which is funded by the European Union. Views and opinions expressed are, however, those of the authors only and do not necessarily reflect those of the European Union, the European Research Executive Agency (REA). Neither the European Union nor the granting authority can be held responsible for them.
Data availability
The data that support the findings of this article are openly available [80].
@@ Appendix A Equation of motion for longitudinal modes
Introducing a canonically normalized mode function $\psi_{L}$ via Eq. (70), we obtain Eq. (71) with the coefficient
@@
For the trivial kinetic coupling, $I_{1}\equiv 1$, this expression coincides with the one reported in Refs. [48, 49]. In the limit $\mu\ll 1$ and for $k\gtrsim aH$, the expression simplifies to the one in Eq. (72). Far inside the horizon, the equation of motion for $\psi_{L}$ in conformal time $\eta$ takes the oscillatorlike form $\psi^{\prime\prime}_{L}(\eta,k)+a^{2}Q\psi_{L}(\eta,k)=0$. To obtain a positive-frequency solution to this equation, we impose the Bunch–Davies boundary condition (79). Using Eq. (70), we translate it to the boundary condition (68) for $\mathcal{A}_{L}$. Further, using Eq. (66), we derive the boundary condition (69) for $\mathcal{D}_{L}$.